OpenStudy (anonymous):

What is the vertex of the graph of y = 2(x – 3)2 + 4?

OpenStudy (anonymous):

this is already written in vertex form: y=a(x-h)^2+k where (h,k) is your vertex. so can you tell me what the vertex to your graph is?

OpenStudy (anonymous):

y = a(x - h)^2 + k y = 2(x – 3)^2 + 4 mach it up. ^^

OpenStudy (anonymous):

these are the options i have (4, 3) (4, –3) (3, 4) (–3, 4)

OpenStudy (anonymous):

what do you mean match it up?